OpenStudy (jack1):

hi guys, quick q, wood love a hand please? if sqrt x = -1 ...does x = 1 fit? or is there no solution to this?

OpenStudy (dumbcow):

impossible!!! sqrt(n) cannot be negative

OpenStudy (jack1):

...hmm but 4 can be (-2)^2 or (+2)^2 ...

OpenStudy (anonymous):

yup sqare root of 1 can be 1 or -1 as

OpenStudy (anonymous):

but square of real number can not be negative

OpenStudy (dumbcow):

ahh \[x^{2} = 4 \rightarrow x = \pm \sqrt{4}\] BUT \[\sqrt{4} = 2 and \sqrt{-4} = impossible\]

OpenStudy (jack1):

gotcha, cheers guys, treat it like a one way street then. Thanks again

OpenStudy (dumbcow):

:)